and, slightly more worringly, here is the probability of throwing a dice and getting at least one six, across 25 attempts - no wonder I'm crap at board games!😁
Any mathematicians out there?
So..it is actually possible to win the top prize more than Once in a year?
What are the odds on a single individual doing that?
So..it is actually possible to win the top prize more than Once in a year?
What are the odds on a single individual doing that?
Approximately a 0.02% chance.
Simon
These odds are surely greater than the weekends lotto draw.😁
What is a good optimal chance of winning?
With 550 prize candidates, at what point (number of potential winners) do the number of prizes have to be increased ( to say 20 prizes ) to maintain fair winning odds?
Compared to most draws the odds are fantastic! I don't think you will get many draws where the top prize is only a 1 in 550 chance when it is worth several thousand pounds and the monthly stake is £2.50. And you get the chance to win 9 other prizes as well! I think we expected rather more people to sign up for this - so make hay whilst you can!
My B&O Icons:
Approximately a 0.02% chance.
Simon
Hmmm - excellent odds, compared to what people otherwise gamble on.
Any one draw = 1 / number of gold members.
Simple solution = remove lots of gold members to increase your odds*......;-)
The chance of winning 1 draw is 1/550
The chance of winning a prize is 120 goes at 1/550.
Then argue if you think this makes you feel you have a better chance of winning anything!
Always wondered why the draw was held in reverse order? if you can win multiple prizes in each draw it doesn't matter what order its done it!
*Now where's my white cat?
Probably of winning in one draw =1/550
Probability of not winning in one draw = 1 - 1/550 = 549/550
Probability of not winning any prizes in 120 draws (a year) = (549/550)^120 = 0.8
Probability of winning more than zero prizes in a year = 1 - Probability of not winning any prizes in 120 draws (a year) = 1 - 0.8 = 0.2 or 20% => 1 in 5
On any particular draw, the probability of winning a prize is still 1 in 550; the draw machine (hopefully) has no memory of previous draws
I did a maths degree over 20 years ago so this may be flakey
Probably of winning in one draw =1/550
Probability of not winning in one draw = 1 - 1/550 = 549/550
Probability of not winning any prizes in 120 draws (a year) = (549/550)^120 = 0.8
Probability of winning more than zero prizes in a year = 1 - Probability of not winning any prizes in 120 draws (a year) = 1 - 0.8 = 0.2 or 20% => 1 in 5
On any particular draw, the probability of winning a prize is still 1 in 550; the draw machine (hopefully) has no memory of previous draws
I did a maths degree over 20 years ago so this may be flakey
(549/550)^120 = 0.804 rather than 0.8000, so you have the same result (80.4% of winning, 19.6% chance of not winning a prize) as a few of us.
Simon
Please Log In to Reply
You must be logged in as a BeoWorld member to post replies to forum threads.
Join BeoWorld
Register to access the forums, reference material, and exclusive member benefits — including our monthly prize draw.
Become a Member